# Deciphering ‘What’ and ‘Where’ Visual Pathways from Spectral Clustering of Layer-Distributed Neural Representations

### v1

@@ Abstract
Regions   Eigs     Image Per-Image Grouping  [Uncaptioned image]
Full Dataset Grouping with Intra- and Inter-Image Affinity   [Uncaptioned image] ‘What’ Pathway [Uncaptioned image] ‘Where’ Pathway
@@
**footnotetext: Equal contribution
@@ 2 Related Work
Spectral clustering of neural features. TokenCut [68], MaskCut [67], and DSM [45] define affinity graphs on top of the final features of a pre-trained DINO [8] model, and use spectral clustering to segment the original image. We take inspiration from these approaches and utilize them as baselines for experimental comparison. Our analysis methodology differs in being global and accounting for features throughout the network, rather than restricted to a single layer. To enable this, we develop a corresponding technical innovation and optimize a more complicated spectral clustering objective via gradient descent, instead of employing a standard eigensolver. This strategy yields improvements on extracting regions from pre-trained DINO. More significantly, our global analysis enables extraction of region structure from models such as CLIP [49] and Stable Diffusion [51], which are not explicitly trained for image correspondence, allowing them to be repurposed for unsupervised semantic segmentation.
@@ 3.1 Spectral Clustering with Distributed Features
$\max_{{\bm{X}}}$ $\mathop{\mathbb{E}}_{{\bm{A}}\in\mathcal{A}}\left[g({\bm{X}})^{\top}{\bm{D}}^{-1}{\bm{A}}g({\bm{X}})\right],$ (3)
@@
where $g(\cdot)$ corresponds to the resampling function that bilinearly interpolates the spatial resolution of ${\bm{X}}$ to match the size of ${\bm{A}}$, allowing affinity matrices to be constructed from feature maps with varying resolutions. This objective encodes a Rayleigh quotient optimization simultaneously across affinities in $\mathcal{A}$, which avoids the intractable exact solution and can naturally scale with increasing $|\mathcal{A}|$.
To optimize, we replace the strict orthogonality requirement with a soft regularization term whose coefficient is $1$. Additionally, notice ${\bm{D}}^{-1}{\bm{A}}$ is a random walk matrix with maximum eigenvalue of $1$. So for numerical stability, we impose a constraint to ensure that the maximum value of the objective does not exceed $1$. Consequently, our final optimization objective can be stated:
$\min_{{\bm{X}}}\mathop{\mathbb{E}}_{{\bm{A}}\in\mathcal{A}}\|g({\bm{X}})^{\top}{\bm{D}}^{-1}{\bm{A}}g({\bm{X}})-1\|_{1}+\lVert{\bm{X}}^{\top}{\bm{X}}-{\bm{I}}\rVert_{2}.$ (5)
We parameterize ${\bm{X}}$ as a learnable feature map and solve for it with gradient-based optimization. In the following subsections, we discuss the choice of affinity set $\mathcal{A}$, and how that choice affects the information we extract.
@@ 3.3 Full-Dataset Extension
With these approximate layer-wise graphs, we optimize the objective in Eqn. 5 a small number of steps per mini-batch, then sample a new mini-batch of images and continue. Finally, this process discovers a consistent set of dense features for a dataset. A visualization of the entire method can be found in Figure 2.
@@ 3.4 Recovering Orthogonal Representations
Eqn. 5 suggests an approximate formulation of the spectral clustering problem. While this results in a structured ${\bm{X}}$, it fails to enforce an orthogonal representation capable of separating distinct features into channels. To overcome this, we orthogonalize ${\bm{X}}$ by finding the eigenvectors ${\bm{U}}$ of a small matrix ${\bm{X}}^{\top}{\bm{X}}\in\mathbb{R}^{C\times C}$. The final representation is given by:
@@ 4.1 Per-Image Region Extraction
Figure 3: Examples of features extracted from different models on PASCAL VOC [21]. Across models we extract meaningful segments. This is true even for models like Stable Diffusion [51], CLIP [49] or MAE [27] whose training is not well-aligned with segmentation.
Figure 4: Oracle-based semantic segmentation performance with varying region count. Across models and number of clusters (regions) returned by K-Means, our method (Ours + K-Means) yields better agreement (in mIoU) with ground-truth than running Normalized Cuts (Ncut + K-Means), or directly applying K-Means on the final output features of the model (K-Means). We observe an even more significant improvement when applying our method to MAE and CLIP, which do not produce discriminative final layer features.
@@
To understand how models partition images spatially, we extract dense eigenvector features for individual images and cluster these features into hard segmentations, as detailed in Section 3.2.
@@
Figure 5: Extracted eigenvectors on COCO for both graph choices. We visualize selected components of ${\bm{X}}_{\text{ortho}}$, sorted by decreasing eigenvalue. Three eigenvectors at a time are rendered as RGB images. In the Q-K case, the first set of eigenvectors describes general scene spatial layout in terms of ground, subject, background, and sky. The second finds top-to-bottom part separation within objects. In the V-V case, the first set of eigenvectors partitions the image into coarse semantics like trees, ground, and sky, while the second set recognizes finer-grained semantic categories and groups individual objects like people, animals, and vehicles.
In Figure 3, we show that features and regions extracted from different models are quite structured, aligning well with object boundaries. We quantify region quality by measuring their oracle overlap with semantic segmentation labels. This gives a sense as to how well attention layers inside models decompose images along semantic axes. We perform this analysis on PASCAL VOC [21], which has 20 foreground classes and 1449 validation images. We score results with the standard metric of mean intersection over union (mIoU) between regions and labels. Each predicted region is assigned to the ground-truth label it overlaps with the most.
@@
To further evaluate the region quality irrespective of decoding choices, we show in Figure 4 the mIoU with varying choice of $K$. We see that the high quality of regions persists across this choice, even when compared with baselines.
Our per-image regions can find broad applicability in a variety of segmentation tasks. For first-step proof-of-concepts see Appendices D.1 and D.2.
@@ 4.2 Full-Dataset Region Extraction
Experimental Setup For constructing graphs, we follow the method in Section 3.3. For efficiency, we concatenate features at each head into a single vector instead of considering heads independently. We select one attention block in the middle block and the first 6 attention blocks in the up-sampling blocks, resulting in a total of 7 attention matrices. We choose channel number $C=50$, cross-image connections $c_{\text{inter}}=c_{\text{intra}}=10$, noise level $t\in\mathcal{U}[20,300)$, and optimize using Adam [33] with a learning rate of 1e-2, and a batch size of 160 images over 4 GPUs for 2100 iterations. When clustering, we choose $K$ to be the number of labels of the relevant task. More details are in Appendix C.
@@ 4.2.1 Qualitative Analysis
By contrast, features from the V-V graph group objects semantically. In COCO, we observe that eigenvectors encode semantic structure hierarchically: the first set of eigenvectors focuses on distinguishing scene-level semantics (e.g., ground, sky, trees) while overlooking differentiating foreground objects. The next set of eigenvectors groups foreground objects like people, animals, and vehicles. In Cityscapes, there is a similar decomposition. The initial set captures broad scene-level semantics, including trees, houses, and the egocentric vehicle, and can differentiate between road and sidewalk. The following set groups cars, people, and road markings. More examples are available in Appendix A.
@@ 4.2.3 Quantitative Analysis for ‘Where’ Pathway
Results on coordinate regression suggest that ${\bm{X}}$ from the Q-K graph contains rich spatial information for processing the ‘where’ pathway. However, both STEGO and the V-V graph group pixels only by semantic similarity and remove spatial information from the final representation. DINO performs well on regression likely due to position embedding information.
@@
Table 4: Results for evaluating ‘where’ pathway on spatial structures. Ours (Q-K graph) outperforms STEGO and ours (V-V graph) in both benchmarks suggesting the Q-K graph contains richer spatial information at object levels. Though DINO can trivially recover the spatial coordinates through positional embeddings, it fails to leverage spatial information for this segmentation task.
@@ 5 Discussion
Our analysis approach is notably distinct from prior work since it requires neither a priori knowledge, nor hyperparameter search over where information is stored in the network. We hope this motivates further work in developing dataset-level analysis tools for peering into the inner workings of deep networks.
@@ D.2 Unsupervised Instance Segmentation
We follow the above procedure on top of the features produced by optimizing Eqn. 5 over Stable Diffusion’s attention layers. We report our results and compare to the current state-of-the-art region proposal methods in Table 7.

### v2

@@ Abstract
Regions   Eigs     Image
Per-Image Grouping
Full Dataset Grouping with Intra- and Inter-Image Affinity
‘What’ Pathway
‘Where’ Pathway
@@
**footnotetext: Equal contribution. Code available at this link.
@@ 2 Related Work
Spectral clustering of neural features. TokenCut [71], MaskCut [70], and DSM [47] define affinity graphs using final features of a pre-trained DINO [8] model, and use spectral clustering to segment the original image. We take inspiration from these approaches and utilize them as baselines for experimental comparison. Our methodology differs in being global and accounting for features throughout the network, rather than restricted to one layer.
@@ 3.1 Spectral Clustering with Distributed Features
$\max_{{\bm{X}}}$ $\mathop{\mathbb{E}}_{{\bm{A}}\in\mathcal{A}}\left[g({\bm{X}})^{\top}{\bm{D}}^{-1}_{{\bm{A}}}{\bm{A}}g({\bm{X}})\right],$
@@
where $g(\cdot)$ corresponds to the resampling function that bilinearly interpolates the spatial resolution of ${\bm{X}}$ to match the size of ${\bm{A}}$, allowing affinity matrices to be constructed from feature maps with varying resolutions. In Eqn. 3.1, we follow Meila and Shi [46] to solve the spectral clustering from the random walk perspective, since the random walk matrix and the attention matrix have the same format and the same eigenvectors. This objective encodes a Rayleigh quotient optimization simultaneously across affinities in $\mathcal{A}$, which avoids the intractable exact solution and can naturally scale with increasing $|\mathcal{A}|$.
Notice that ${\bm{D}}^{-1}_{{\bm{A}}}{\bm{A}}$ is a random walk matrix with maximum eigenvalue of $1$. For numerical stability, we impose a constraint to ensure that the maximum value of the objective does not exceed $1$. In addition, we replace the strict orthogonality requirement with a soft Frobenius regularization term whose coefficient is $1$. Consequently, our final optimization objective is:
$\min_{{\bm{X}}}\mathop{\mathbb{E}}_{{\bm{A}}\in\mathcal{A}}|g({\bm{X}})^{\top}{\bm{D}}^{-1}_{{\bm{A}}}{\bm{A}}g({\bm{X}})-1|+\lVert{\bm{X}}^{\top}{\bm{X}}-{\bm{I}}\rVert_{F}.$ (4)
We parameterize ${\bm{X}}$ as a learnable feature map and solve for it using gradient-based optimization. In the following subsections, we discuss the choice of affinity set $\mathcal{A}$ and how that choice affects the information we extract.
@@ 3.3 Full-Dataset Extension
With these approximate layer-wise graphs, we optimize the objective in Eqn. 4 a small number of steps per mini-batch, then sample a new mini-batch of images and continue. Finally, this process discovers a consistent set of dense features for a dataset. A visualization of the entire method can be found in Figure 2.
@@ 3.4 Recovering Orthogonal Representations
Eqn. 4 suggests an approximate formulation of the spectral clustering problem. While this results in a structured ${\bm{X}}$, it fails to enforce an orthogonal representation capable of separating distinct features into channels. To overcome this, we orthogonalize ${\bm{X}}$ by finding the eigenvectors ${\bm{U}}$ of a small matrix ${\bm{X}}^{\top}{\bm{X}}\in\mathbb{R}^{C\times C}$. This is similar to the reorthogonalization step in approximate eigensolvers; e.g., lines 36-38 of Algorithm 2 in Maire and Yu [42]. The final representation is given by:
@@ 4.1 Per-Image Region Extraction
Figure 3: Features extracted from different models on PASCAL VOC [21]. Across models we extract meaningful regions, even for models like Stable Diffusion [53], CLIP [51] or MAE [27] whose training is not well-aligned with segmentation.
Figure 4: Oracle-based semantic segmentation performance with varying region count. Across models and number of clusters (regions) returned by K-Means, our method (Ours + K-Means) yields better agreement (in mIoU) with ground-truth than running Normalized Cuts (Ncut + K-Means), or directly applying K-Means on the final output features of the model (K-Means). We observe an even more significant improvement when applying our method to MAE and CLIP, which do not produce discriminative features.
@@
To show how models partition images spatially, we extract dense eigenvector for individual images and cluster these features into hard segmentations, as detailed in Section 3.2.
Config All Enc Mid Dec 32x32 64x64
 Layer Index 1-16 1-4 5 - 10 11-16
3-4
11-13
1-2
14-16
mIoU 0.75 0.66 0.76 0.80 0.75 0.70
$t_{\text{max}}$ 250 500 750 999
  mIoU 0.77 0.75 0.74 0.69
Table 2: Ablation of layer index and maximum noise level of the diffusion model on the PASCAL VOC dataset [21]. We find that using only decoder layers and middle noise yields the best results.
@@
Figure 5: Extracted eigenvectors on COCO for both graph choices. We visualize selected components of ${\bm{X}}_{\text{ortho}}$, sorted by decreasing eigenvalue. Three eigenvectors at a time are rendered as RGB images. In the Q-K case, the first set of eigenvectors describes general scene spatial layout in terms of ground, subject, background, and sky. The second finds top-to-bottom part separation within objects. In the V-V case, the first set of eigenvectors partitions the image into coarse semantics like trees, ground, and sky, while the second set recognizes finer-grained categories and groups individual objects like people, animals, and vehicles.
In Figure 3, we show that features and regions extracted from different models are quite structured, aligning well with object boundaries. We quantify region quality by measuring their oracle overlap with semantic segmentation labels. This gives a sense as to how well attention layers inside models decompose images along semantic axes. We perform this analysis on PASCAL VOC [21], which has 20 foreground classes and 1449 validation images. We score results with the metric of mean intersection over union (mIoU) between regions and labels. Each region is assigned to the ground-truth label it overlaps with the most.
@@
In Table 2, we provide ablation studies on the choices of layer for feature extraction and maximum noise level. Stable diffusion has 16 attention layers with resolutions from 64x64 to 8x8. Our default is All (1-16), $t_{\text{max}}$ = 500. We conduct experiments on a per-image region extraction setting with 200 images from the PASCAL VOC validation set. Although our main experiment utilizes features from all layers, making minimal assumptions about layer-wise feature distribution, we find that using only decoder layers and a middle noise level yields better results.
To further evaluate the region quality irrespective of decoding choices, Figure 4 shows the mIoU with varying choice of $K$. We see that the high quality of regions persists across choices, even when compared with baselines.
Our per-image regions can find broad applicability in a variety of segmentation tasks. For first-step proof-of-concepts, see Appendices D.1 and D.2.
@@ 4.2 Full-Dataset Region Extraction
Experimental Setup. For constructing graphs, we follow the method in Section 3.3. For efficiency, we concatenate features at each head into a single vector instead of considering heads independently. We select one attention block in the middle block and the first 6 attention blocks in the up-sampling blocks, resulting in a total of 7 attention matrices. We choose channel number $C=50$, cross-image connections $c_{\text{inter}}=c_{\text{intra}}=10$, noise level $t\in\mathcal{U}[20,300)$, and optimize using Adam [33] with a learning rate of 1e-2, and a batch size of 160 images over 4 GPUs for 2100 iterations. When clustering, we choose $K$ to be the number of labels of the relevant task. More details are in Appendix C.
@@ 4.2.1 Qualitative Analysis
By contrast, features from the V-V graph group objects semantically. In COCO, we observe that eigenvectors encode semantic structure hierarchically: the first set of eigenvectors focuses on distinguishing scene-level semantics (e.g., ground, sky, trees) while overlooking differentiating foreground objects. The next set of eigenvectors groups foreground objects like people, animals, and vehicles. In Cityscapes, the initial set captures broad scene-level semantics, including trees, houses, and the egocentric vehicle, and can differentiate between road and sidewalk. The following set groups cars, people, and road markings. More examples are available in Appendix A.
@@ 4.2.3 Quantitative Analysis for ‘Where’ Pathway
Results on coordinate regression suggest that ${\bm{X}}$ from the Q-K graph contains rich spatial information for processing the ‘where’ pathway. However, both STEGO and the V-V graph group pixels only by semantic similarity and remove spatial information from the final representation. DINO performs well on regression likely due to position embeddings.
@@
Table 5: Results for evaluating ‘where’ pathway on spatial structures. Ours (Q-K graph) outperforms STEGO and ours (V-V graph) in both benchmarks suggesting the Q-K graph contains richer spatial information at object levels. Though DINO can trivially recover the spatial coordinates through positional embeddings, it fails to leverage that information for segmentation.
@@ 5 Discussion
Our approach could be the first example in a new class of optimization-centric techniques for peering into the inner workings of deep networks. Future research could repurpose other computationally intensive, but scalable, classic machine learning tools to the analysis of network behavior.
Acknowledgments. NSF GRFP (1754881) supported D. Yunis.
@@ D.2 Unsupervised Instance Segmentation
We follow the above procedure on top of the features produced by optimizing Eqn. 4 over Stable Diffusion’s attention layers. We report our results and compare to the current state-of-the-art region proposal methods in Table 8.